(3) A survey revealed that the mean cholesterol level of male athletes in Ghana aged 20 – 30 years is 180 mg/dl and the standard deviation is 43 mg/dl. If a random sample of 60 male athletes is selected from this population, what is the probability that the sample mean cholesterol level will be greater than 190 mg/dl?
1
Expert's answer
2020-07-15T19:14:56-0400

given : mean = 180

Sd= 43, n= 60


to find :

P("\overline{X}" > 190)


solution :


P("\overline{X}" > 190)= P("\frac{\overline{X}-mean}{\frac{sd}{\sqrt{n}}}" > "\frac{190-180}{\frac{43}{\sqrt{60}}}" )

=P(Z>1.80)

=1- P(Z<1.80)

Using z table , 1- 0.9642

= 0.0358